# Efficient Individually Rational Recommender System under Stochastic Order

### v1

@@ Appendix G Proofs of Observations and Propositions from Sections 3 and 4
    hence, $\boldsymbol{p}_{i,j}$ is IR w.r.t. to $\mathcal{I}$. After selecting $\boldsymbol{p}_{i,j}$, either $a_{i}$ was realized or $a_{j}$. In the former, the information remains the same, and we can repeat this experiment again. The probability of $a_{j}$ realizing is positive and constant, and hence, after a finite time, we will eventually realize it. Once we do, the number of unexplored armed decreases by one. We can follow this process until all arms are explored. ∎
  • Proof of Observation 4.1.The proof of this observation relies on constructing a policy $\pi$ that simulates ALG. Since by definition $W(\pi,A)\leq W^{\star}(A)$, it is enough to show that $\lim_{T\rightarrow\infty}SW_{T}({\small\textnormal{{ALG}}})\leq W(\pi,A)$. In every round, $\pi$ selects precisely what ALG selects, and if the realized arm was already explored by ALG, $\pi$ ignores it. The infinite time expected value of ALG cannot exceed $W(\pi,A)$. The full details are similar to [5, Theorem 3] and are hence omitted. ∎
  • Proof of Observation 4.2.Fix any policy $\pi$. Let ${\small\textnormal{{ALG}}}(\pi)$ be the modification of Algorithm 2 that uses $\pi$ instead of OGP in Lines 2-2. Once $\pi$ reaches a terminal state, ${\small\textnormal{{ALG}}}(\pi)$ secures the reward of $\pi$ in finite time. Overall, $\lim_{T\rightarrow\infty}SW_{T}({\small\textnormal{{ALG}}}(\pi))=W(\pi,A)$. ∎
  • Proof of Proposition 4.1.Fix an IRSR instance such that $(X(a_{i}))_{i}\in\{x^{-},x^{+}\}$ (for $x^{-}\leq x^{+}$) almost surely. For the problem to be non-trivial, we must have $x^{-}<0$ and $x^{+}>0$. Otherwise, if $x^{-},x^{+}<0$ the only IR action is $a_{0}$, and if $x^{-},x^{+}\geq 0$, we can explore all arms using the singleton portfolios $(\boldsymbol{p}_{ii})_{i\in[K]}$. From here on, we assume w.l.o.g. that $x^{-}=-1$ and $x^{+}=H$. For convenience, we state ${\small\textnormal{{IREGB}}}^{\prime}$ explicitly in Algorithm 4. Before we prove the proposition, we remark that
@@
      Since $(X(a_{i}))_{i}$ take either $-1$ or $H$, Assumption 2.1 implies a stochastic order on all arms, not only on $\textnormal{neg}(A)$.
@@
      This proof uses the analysis presented in Section A.
@@
Step 1: In the case of realizing a positive reward, any algorithm would stop exploring and exploit that realized reward. Consequently, we can separate exploration rounds from exploitation rounds. Notice, however, that the exploration policy can select portfolios different that OGP for finite $T$. To illustrate, reconsider Example A.1. In the extreme case of $T=1$, there is no point in selecting $\boldsymbol{p}_{1,3}$, since exploring $a_{3}$ is futile; we only care about maximizing the current round’s reward.
@@
Using similar arguments to those in Proposition A.1, we can assume w.l.o.g. that $\pi$ belongs to $\{2^{A}\rightarrow\mathcal{P}\cup\mathcal{P}^{\prime}\}$ (recall the definition of $\mathcal{P}$ and $\mathcal{P}^{\prime}$ from Subsection A.2). To see this, observe that any IR policy can be formulated as a convex combination of policies that use $\mathcal{P}\cup\mathcal{P}^{\prime}$ solely, and therefore we can assume that $\pi$ is the one for which ${\small\textnormal{{ALG}}}^{\pi}$ gets the highest social welfare. Furthermore, due to the proof of Theorem A.1 (precisely Equation (E)) it follows that if $\pi$ does not admit the structure of OGP, then it is strictly sub-optimal. Next, let
@@
Step 2: Notice, however, that OGP is a class of policies differing from one another in the choices of arms from $\textnormal{pos}(A)$ (Line 1 in Policy 1); hence, one policy may attain a better reward than the other by reaching exploitation faster.
As we commented in Line 4 of Algorithm 4, when the state $s$ does not contain arms from $\textnormal{neg}(A)$ we prioritize singleton portfolios according to the stochastic order. That is, we favor $\boldsymbol{p}_{i,i}$ over $\boldsymbol{p}_{i^{\prime},i^{\prime}}$ if $\mu(a_{i})>\mu(a_{i^{\prime}})$. This modification ensures that the time to exploitation from such states is minimal.
@@
Input: the OGP policy
1 $s\leftarrow A$
2 while ${\small\textnormal{{OGP}}}(s)\neq\emptyset$ do
      // $s$ is not a terminal state
3      select ${\small\textnormal{{OGP}}}(s)$, and denote the realized action by $a_{k}$
       // if $\textnormal{neg}(s)=\emptyset$, prioritize the arms in $\textnormal{pos}(s)$ according to the stochastic order
4      if $x_{a_{k}}>0$ then
            // a reward of $H$ was realized
5            break
6
7      $s\leftarrow s\setminus\{a_{k}\}$
8
9if $x(a_{k})=H$ for some explored arm $a_{k}$ then
10      exploit $a_{k}$ forever
11
12else
13      exploit $a_{0}$ forever
14
@@ 7      $s\leftarrow s\setminus\{a_{k}\}$
Since all OGP have the same expected exploration time, they all achieve the same social welfare. This completes the proof of Proposition 4.1. ∎
  • Proof of Proposition 4.2.To prove the claim, we show that IREGB explores for at most $K(1+\frac{\eta}{\delta})$ rounds, and then exploits. First, IREGB uses $K_{1}\leq K$ rounds following OGP until it reaches a terminal state (the while loop in Line 2 breaks). If all the realized rewards are negative, the exploration ends after that. Otherwise, if it discovers a positive reward, the value of that reward is at least $\delta$. Next, the Bernoulli trails will explore every unexplored arm w.p. of at least $\frac{\delta}{\delta+\eta}$, or $\frac{\delta+\eta}{\delta}$ rounds in expectation. Therefore, after $(K-K_{1})(1+\frac{\eta}{\delta})$ rounds in expectation we explore all the remaining arms. Overall, the expected number of rounds devoted to exploration is
@@
    This completes the proof of Proposition 4.2. ∎
Appendix H Proof of Statements from Section A
  • Proof of Proposition A.1.Fix a non-terminal state $U\subseteq A$. Further, for simplicity of notation, denote $V(a)\stackrel{{\scriptstyle\text{def}}}{{=}}W^{*}(U\setminus\{a\})$ for every $a\in U$. Due to Equation (7), the action that maximizes the reward at state $U$ is the solution $\boldsymbol{p}\in\Delta(U)$ of the following linear program:
@@ Appendix H Proof of Statements from Section A
    Finally, notice that the constraint $\sum_{i,j}\alpha_{i,j}\left(\boldsymbol{p}_{i,j}(i)\mu(a_{i})+\boldsymbol{p}_{i,j}(j)\mu(a_{j})\right)\geq 0$ holds for every selection of $(\alpha_{i,j})_{i,j}$ by the way we defined $\mathcal{P}\cup\mathcal{P}^{\prime}$; thus, the maximum of Problem (P2) is obtained when we set $\alpha_{i,j}=1$ for the pair $(i,j)$ that maximizes $\left(\boldsymbol{p}_{i,j}(i)V(a_{i})+\boldsymbol{p}_{i,j}(j)V(a_{j})\right)$. ∎
Claim 5.
Consider a state $U\in\mathcal{S}$, such that $\textnormal{neg}(U)\geq 2$. Let $a_{j}=\operatorname*{arg\,min}_{a_{j^{\prime}}\in\textnormal{neg}(U)}\sigma^{\textnormal{right}}_{\pi}(a_{j^{\prime}})$, and let $a_{\tilde{j}}\in\textnormal{neg}(U),a_{\tilde{j}}\neq a_{j}$. Under Assumption 2.1, it might be the case that $W^{*}(U\setminus\{a_{j}\})<W^{*}(U\setminus\{a_{\tilde{j}}\})$.
  • Proof of Claim 5.We prove the claim by providing an example, that could be easily extended to a family of infinitely many examples. Consider $K=3$, $A=\{a_{1},a_{2},a_{3}\}$ such that
@@
    The proof is completed by taking $\epsilon$ to zero. ∎
Appendix I Bayesian Incentive Compatibility
  • Proof of Proposition 5.1.The proof follows closely the proof of Bahar et al. [5, Section E]; hence, we omit the details. ∎
  • Proof of Theorem 5.1.IC-IREGB is IR by design, as it only recommends the default arm, a greedy arm, or according to IREGB. To analyze its social welfare, observe that after $K(1+{\eta}\operatorname{\mathbb{E}}\left[\frac{1}{\delta}\right])$ phases in expectation, we only explore. This follows directly from Proposition 4.2. Since the length of each phase is $B=\left\lceil\frac{H}{\xi\gamma}\right\rceil+1$, there are at most $O\left(\frac{K\eta H\operatorname{\mathbb{E}}\left[\frac{1}{\delta}\right]}{\xi\gamma}\right)$ exploration rounds in expectation. To show it is BIC, we divide the analysis according to the agent’s index. The analysis closely follows the proof of Bahar et al. [5, Theorem 3], and is hence omitted. ∎

### v2

@@ Appendix 0.G Proofs of Observations and Propositions from Sections 3 and 4
hence, $\boldsymbol{p}_{i,j}$ is MIR w.r.t. to $\mathcal{I}$. After selecting $\boldsymbol{p}_{i,j}$, either $a_{i}$ was realized or $a_{j}$. In the former, the information remains the same, and we can repeat this experiment again. The probability of $a_{j}$ realizing is positive and constant, and hence, after a finite time, we will eventually realize it. Once we do, the number of unexplored armed decreases by one. We can follow this process until all arms are explored.
Proof (Proof of Observation 2)
The proof of this observation relies on constructing a policy $\pi$ that simulates ALG. Since by definition $W(\pi,A)\leq W^{\star}(A)$, it is enough to show that $\lim_{T\rightarrow\infty}SW_{T}({\small\textnormal{{ALG}}})\leq W(\pi,A)$. In every round, $\pi$ selects precisely what ALG selects, and if the realized arm was already explored by ALG, $\pi$ ignores it. The infinite time expected value of ALG cannot exceed $W(\pi,A)$. The full details are similar to [6, Theorem 3] and are hence omitted.
Proof (Proof of Observation 3)
Fix any policy $\pi$. Let ${\small\textnormal{{ALG}}}(\pi)$ be the modification of Algorithm 1 that uses $\pi$ instead of OGP in Lines 2-3. Once $\pi$ reaches a terminal state, ${\small\textnormal{{ALG}}}(\pi)$ secures the reward of $\pi$ in finite time. Overall, $\lim_{T\rightarrow\infty}SW_{T}({\small\textnormal{{ALG}}}(\pi))=W(\pi,A)$.
Proof (Proof of Proposition 2)
Fix an IRSR instance such that $(X(a_{i}))_{i}\in\{x^{-},x^{+}\}$ (for $x^{-}\leq x^{+}$) almost surely. For the problem to be non-trivial, we must have $x^{-}<0$ and $x^{+}>0$. Otherwise, if $x^{-},x^{+}<0$ the only MIR action is $a_{0}$, and if $x^{-},x^{+}\geq 0$, we can explore all arms using the singleton portfolios $(\boldsymbol{p}_{ii})_{i\in[K]}$. From here on, we assume w.l.o.g. that $x^{-}=-1$ and $x^{+}=H$. For convenience, we state ${\small\textnormal{{IREGB}}}^{\prime}$ explicitly in Algorithm 3. Before we prove the proposition, we remark that
@@
    Since $(X(a_{i}))_{i}$ take either $-1$ or $H$, Assumption 1 implies a stochastic order on all arms, not only on $\textnormal{neg}(A)$.
@@
    This proof uses the analysis presented in Section 0.A.
@@
Step 1: In the case of realizing a positive reward, any algorithm would stop exploring and exploit that realized reward. Consequently, we can separate exploration rounds from exploitation rounds. Notice, however, that the exploration policy can select portfolios different that OGP for finite $T$. To illustrate, reconsider Example 2. In the extreme case of $T=1$, there is no point in selecting $\boldsymbol{p}_{1,3}$, since exploring $a_{3}$ is futile; we only care about maximizing the current round’s reward.
@@
Using similar arguments to those in Proposition 5, we can assume w.l.o.g. that $\pi$ belongs to $\{2^{A}\rightarrow\mathcal{P}\cup\mathcal{P}^{\prime}\}$ (recall the definition of $\mathcal{P}$ and $\mathcal{P}^{\prime}$ from Subsection 0.A.2). To see this, observe that any MIR policy can be formulated as a convex combination of policies that use $\mathcal{P}\cup\mathcal{P}^{\prime}$ solely, and therefore we can assume that $\pi$ is the one for which ${\small\textnormal{{ALG}}}^{\pi}$ gets the highest social welfare. Furthermore, due to the proof of Theorem 0.A.1 (precisely Equation (0.E)) it follows that if $\pi$ does not admit the structure of OGP, then it is strictly sub-optimal. Next, let
@@
Step 2: Notice, however, that OGP is a class of policies differing from one another in the choices of arms from $\textnormal{pos}(A)$ (Line 4 in Policy 1); hence, one policy may attain a better reward than the other by reaching exploitation faster.
As we commented in Line 3 of Algorithm 3, when the state $s$ does not contain arms from $\textnormal{neg}(A)$ we prioritize singleton portfolios according to the stochastic order. That is, we favor $\boldsymbol{p}_{i,i}$ over $\boldsymbol{p}_{i^{\prime},i^{\prime}}$ if $\mu(a_{i})>\mu(a_{i^{\prime}})$. This modification ensures that the time to exploitation from such states is minimal.
@@
0:  the OGP policy
1:  $s\leftarrow A$
2:  while ${\small\textnormal{{OGP}}}(s)\neq\emptyset$ do #$s$ is not a terminal state
3:     select ${\small\textnormal{{OGP}}}(s)$, and denote the realized action by $a_{k}$if $\textnormal{neg}(s)=\emptyset$, prioritize the arms in $\textnormal{pos}(s)$ according to the stochastic order
4:     if $x_{a_{k}}>0$ then #a reward of $H$ was realized
5:        break
6:     $s\leftarrow s\setminus\{a_{k}\}$
7:  if $x(a_{k})=H$ for some explored arm $a_{k}$ then
8:     exploit $a_{k}$ forever
9:  else
10:     exploit $a_{0}$ forever
Since all OGP have the same expected exploration time, they all achieve the same social welfare. This completes the proof of Proposition 2.
Proof (Proof of Proposition 3)
To prove the claim, we show that IREGB explores for at most $K(1+\frac{\eta}{\delta})$ rounds, and then exploits. First, IREGB uses $K_{1}\leq K$ rounds following OGP until it reaches a terminal state (the while loop in Line 2 breaks). If all the realized rewards are negative, the exploration ends after that. Otherwise, if it discovers a positive reward, the value of that reward is at least $\delta$. Next, the Bernoulli trials will explore every unexplored arm w.p. of at least $\frac{\delta}{\delta+\eta}$, or $\frac{\delta+\eta}{\delta}$ rounds in expectation. Therefore, after $(K-K_{1})(1+\frac{\eta}{\delta})$ rounds in expectation we explore all the remaining arms. Overall, the expected number of rounds devoted to exploration is
@@
This completes the proof of Proposition 3.
Appendix 0.H Proof of Statements from Section 0.A
Proof (Proof of Proposition 5)
Fix a non-terminal state $U\subseteq A$. Further, for simplicity of notation, denote $V(a)\stackrel{{\scriptstyle\text{def}}}{{=}}W^{*}(U\setminus\{a\})$ for every $a\in U$. Due to Equation (7), the action that maximizes the reward at state $U$ is the solution $\boldsymbol{p}\in\Delta(U)$ of the following linear program:
@@
Finally, notice that the constraint $\sum_{i,j}\alpha_{i,j}\left(\boldsymbol{p}_{i,j}(i)\mu(a_{i})+\boldsymbol{p}_{i,j}(j)\mu(a_{j})\right)\geq 0$ holds for every selection of $(\alpha_{i,j})_{i,j}$ by the way we defined $\mathcal{P}\cup\mathcal{P}^{\prime}$; thus, the maximum of Problem (P2) is obtained when we set $\alpha_{i,j}=1$ for the pair $(i,j)$ that maximizes $\left(\boldsymbol{p}_{i,j}(i)V(a_{i})+\boldsymbol{p}_{i,j}(j)V(a_{j})\right)$.
Claim 5
Consider a state $U\in\mathcal{S}$, such that $\textnormal{neg}(U)\geq 2$. Let $a_{j}=\operatorname*{arg\,min}_{a_{j^{\prime}}\in\textnormal{neg}(U)}\sigma^{\textnormal{right}}_{\pi}(a_{j^{\prime}})$, and let $a_{\tilde{j}}\in\textnormal{neg}(U),a_{\tilde{j}}\neq a_{j}$. Under Assumption 1, it might be the case that $W^{*}(U\setminus\{a_{j}\})<W^{*}(U\setminus\{a_{\tilde{j}}\})$.
Proof (Proof of Claim 5)
We prove the claim by providing an example, that could be easily extended to a family of infinitely many examples. Consider $K=3$, $A=\{a_{1},a_{2},a_{3}\}$ such that
@@
The proof is completed by taking $\epsilon$ to zero.
Appendix 0.I Bayesian Incentive Compatibility
Proof (Proof of Proposition 4)
The proof follows closely the proof of Bahar et al. [6, Section E]; hence, we omit the details.
Proof (Proof of Theorem 5.1)
BIC-IREGB is MIR by design, as it only recommends the default arm, a greedy arm, or according to IREGB. To analyze its social welfare, observe that after $K(1+{\eta}\operatorname{\mathbb{E}}\left[\frac{1}{\delta}\right])$ phases in expectation, we only explore. This follows directly from Proposition 3. Since the length of each phase is $B=\left\lceil\frac{H}{\xi\gamma}\right\rceil+1$, there are at most $O\left(\frac{K\eta H\operatorname{\mathbb{E}}\left[\frac{1}{\delta}\right]}{\xi\gamma}\right)$ exploration rounds in expectation. To show it is BIC, we divide the analysis according to the agent’s index. The analysis closely follows the proof of Bahar et al. [6, Theorem 3], and is hence omitted.